To design a new advertising campaign, Ford Motor Company would like to estimate the proportion of drivers of the new Ford Fusion that are women. In a random sample of 91 Fusion owners, 49 of them were women. What is the 99% confidence interval estimating the proportion of all drivers who are women?

Question options:

1)

( 0.32693 , 0.59615 )

2)

( 0.41689 , 0.66003 )

3)

( -0.40385 , 0.67307 )

4)

( 0.40385 , 0.67307 )

5)

( 0.4862 , 0.59072 )

Question (1 point)

The owner of a local supermarket believes the average number of gallons of milk the store sells per day is 197.3. In a random sample of 25 days, the owner finds that the average number of gallons sold was 205.9 with a standard deviation of 37.28. Using this information, the owner calculated the confidence interval of (190.5, 221.3) with a confidence level of 95%. Which of the following statements is the best conclusion?

Question options:

1)

We are 95% confident that the average number of gallons sold per day is less than 197.3.

2)

The percentage of days on which more than 197.3 gallons of milk are sold is 95%.

3)

We cannot determine the proper interpretation based on the information given.

4)

We are 95% confident that the average number of gallons sold per day is greater than 197.3.

5)

The average number of gallons sold per day is not signficantly different from 197.3.
Answer #1

1)

4th option

Standard error of the mean = SEM = √x(N-x)/N3 = 0.052

α = (1-CL)/2 = 0.005

Standard normal deviate for α = Zα = 2.576

Proportion of positive results = P = x/N = 0.538

Lower bound = P - (Zα*SEM) = 0.404

Upper bound = P + (Zα*SEM) = 0.673

2)

5)

The average number of gallons sold per day is not signficantly different from 197.3.